OpenStudy (anonymous):

In her first five basketball games, Tara scored 10, 11, 15, 16, and 18 points. What is the mean absolute deviation for this set of scores? I think the answer is 14. Let me know if I'm right if not can someone help me find out what the right answer is?

OpenStudy (anonymous):

okay the mean is all the numbers added together and then divide by how many number there is. for your case: mean = 10+11+15+16+18=70 there is 5 numbers so... 70/5= 14 which you are correct. Hope this helps!

OpenStudy (anonymous):

yup that's how I did it I just wanted to know if it was right Thanks!

OpenStudy (anonymous):

your welcome